OpenStudy (anonymous):

What is the volume of the cube? A. 24 cubic inches B. 96 cubic inches C. 384 cubic inches D. 512 cubic inches

OpenStudy (anonymous):

what's the cube look like

OpenStudy (anonymous):

THE ANSWER IS.... D

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

yes I am sure

OpenStudy (anonymous):

8 x 8 x 8 so yeah gswag98 is correct, a^3 = cube :) 512 is the answer!

OpenStudy (anonymous):

are you gonna give me a medal @Roc07or

OpenStudy (anonymous):

first you do 8 x 8 for the face of the cube (which is 64) then you count how much faces the cube has (which it has 6 faces) then tou do 64 x 6 which =384, you don't just do the numbers on the cube,the numbers cube are just an example of how much are on each face so the answer would be c not d

OpenStudy (triciaal):

volume of a cube = length *width*height and all these are the same number of units volume = L^3 for this problem l=8 therefore volume = 8^3 = 512